Makers Corp. had additions to retained earnings for the year just ended of $328,000. The firm paid out $176,000 in cash dividends, and it has ending total equity of $4.81 million. The company currently has 140,000 shares of common stock outstanding. a. What are earnings per share? (Do not round intermediate calculations and round b. What are dividends per share? (Do not round intermediate calculations and round c. What is the book value per share? (Do not round intermediate calculations and d. If the stock currently sells for $60 per share, what is the market-to-book ratio? (Do your answer to 2 decimal places, e.g., 32.16.) your answer to 2 decimal places, e.g., 32.16.) round your answer to 2 decimal places, e.g., 32.16.) not round intermediate calculations and round your answer to 2 decimal places, e.g, 32.16.) your answer to 2 decimal places, e.g., 32.16.) Intermediate calculations and round your answer to 2 decimal places, e.g, 32.16.) e. What is the price-earnings ratio? (Do not round intermediate calculations and round f If the company had sales of $4.55 million, what is the price-sales ratio? (Do not round a. Earnings per share b. Dividends per share c. Book value per share d. Market-to-book ratio e Price-earnings ratio f. Price-sales ratio times times times <Prev 4of 101E1 Nex 25 MacBook Air
Answer #1
a. EPS = Net Income / Common Stock outstanding
EPS = (Addition to RE 328000+ Cash dividend 176000) / Common Stock outstanding 140000
EPS = $3.60 per Common Share
b. Dividend per share = dividend amount / Common Stock outstanding
Dividend per share = 176000 / 140000 = $1.26 per Common share
c. Book Value per share = Total Equity at end year / Common Stock outstanding
Book Value per share = 4810000/ 140000 = $34.36 per Common share
d. Market to book ratio = Market price / Book value = $60 / $34.36 = 1.75 times
e. Price to Earnings ratio = Market price / EPS = $60 / $3.60 = 16.67 times
f. Price to Sales ratio = Total market value of equity / Total sales of company =
140000* $60 / $4550000 = 0.54 times